OpenStudy (anonymous):

What is the percent decrease from 60 to 39? no choices

OpenStudy (anonymous):

http://brainly.com/question/286832 go here

OpenStudy (anonymous):

1) 100 / 60 = (5/3) 2) (5/3) * 39 = 65% 3) 100% - 65% = 35% decrease

OpenStudy (anonymous):

ok sorry i went to go eat

Directrix (directrix):

To calculate percentage decrease: First: work out the difference (decrease) between the two numbers you are comparing. Then: divide the decrease by the original number and multiply the answer by 100. If your answer is a negative number then this is a percentage increase.